# ReconX: Reconstruct Any Scene from Sparse Views with Video Diffusion Model

### v1

@@ Abstract
Advancements in 3D scene reconstruction have transformed 2D images from the real world into 3D models, producing realistic 3D results from hundreds of input photos. Despite great success in dense-view reconstruction scenarios, rendering a detailed scene from insufficient captured views is still an ill-posed optimization problem, often resulting in artifacts and distortions in unseen areas. In this paper, we propose ReconX, a novel 3D scene reconstruction paradigm that reframes the ambiguous reconstruction challenge as a temporal generation task. The key insight is to unleash the strong generative prior of large pre-trained video diffusion models for sparse-view reconstruction. However, 3D view consistency struggles to be accurately preserved in directly generated video frames from pre-trained models. To address this, given limited input views, the proposed ReconX first constructs a global point cloud and encodes it into a contextual space as the 3D structure condition. Guided by the condition, the video diffusion model then synthesizes video frames that are both detail-preserved and exhibit a high degree of 3D consistency, ensuring the coherence of the scene from various perspectives. Finally, we recover the 3D scene from the generated video through a confidence-aware 3D Gaussian Splatting optimization scheme. Extensive experiments on various real-world datasets show the superiority of our ReconX over state-of-the-art methods in terms of quality and generalizability. Project Page: https://liuff19.github.io/ReconX.
@@ 1 Introduction
To address the limitations, we propose ReconX, a novel 3D scene reconstruction paradigm that reformulates the inherently ambiguous reconstruction problem as a generation problem. Our key insight is to unleash the strong generative prior of pre-trained large video diffusion models (Blattmann et al., 2023a; b; Xing et al., 2023) to create more observations for the downstream reconstruction task. Despite the capability to synthesize video clips featuring plausible 3D structure (Gao et al., 2024), recovering a high-quality 3D scene from current video diffusion models is still challenging due to the poor 3D view consistency across generated 2D frames. Grounded by theoretical analysis, we explore the potential of incorporating 3D structure guidance into the video generative process, which bridges the gap between the under-determined 3D creation problem and the fully-observed 3D reconstruction setting. Specifically, given sparse images, we first build a global point cloud through a pose-free stereo reconstruction method. Then we encode it into a rich context representation space as the 3D condition in cross-attention layers, which guides the video diffusion model to synthesize detail-preserved frames with 3D consistent novel observations of the scene. Finally, we reconstruct the 3D scene from the generated video through Gaussian Splatting with a 3D confidence-aware and robust scene optimization scheme, which further deblurs the uncertainty in video frames effectively. Extensive experiments verify the efficacy of our framework and show that ReconX outperforms existing methods for high quality and generalizability, revealing the great potential to craft intricate 3D worlds from video diffusion models. The overview and examples of reconstructions are shown in Figure 1.
@@
In summary, our main contributions are:
@@
  • We incorporate the 3D structure guidance into the conditional space of the video diffusion model to generate 3D consistent frames and propose a 3D confidence-aware optimization scheme in 3DGS to reconstruct the scene given the generated video.
@@ 2 Related Work
Sparse-view reconstruction. NeRF and 3DGS typically demand hundreds of input images and rely on the multi-view stereo reconstruction (MVS) approaches (e.g., COLMAP (Schönberger & Frahm, 2016)) to estimate the camera parameters. To address the issue of low-quality 3D reconstruction caused by sparse views, PixelNeRF (Yu et al., 2021) proposes using convolutional neural networks to extract features from the input context. Moreover, FreeNeRF (Yang et al., 2023) adopts the frequency and density regularized strategies to alleviate the artifacts caused by insufficient inputs without any additional cost. To mitigate the overfitting to input sparse views in 3DGS, FSGS (Zhu et al., 2023) and SparseGS (Xiong et al., 2023) employ a depth estimator to regularize the optimization process. However, these methods all require known camera intrinsics and extrinsics, which is not practical in real-world scenario. Benefiting from the existing powerful 3D reconstruction model (i.e., DUSt3R (Wang et al., 2024a)), InstantSplat (Fan et al., 2024) is able to acquire accurate camera parameters and initial 3D representations from unposed sparse-view inputs, leading to the efficient and high-quality 3D reconstruction.
@@
Regression model for generalizable view synthesis. While NeRF and 3DGS are optimized per-scene, a line of research aims to train feed-forward models that output a 3D representation directly from a few input images, bypassing the need for time-consuming optimization. Splatter image (Szymanowicz et al., 2024b) performs an efficient feed-forward manner for monocular 3D object reconstruction by predicting a 3D Gaussian for each image pixel. Meanwhile, pixelSplat (Charatan et al., 2024) proposes predicting the scene-level 3DGS from the image pairs, using the epipolar transformer to better extract scene features. Following that, MVsplat (Chen et al., 2024a) introduces the cost volume and depth refinements to produce a clean and high-quality 3D Gaussians in a faster way. To reconstruct a complete scene from a single image, Flash3D (Szymanowicz et al., 2024a) adopts a hierarchical 3DGS learning policy and depth constraint to achieve high-quality interpolation and extrapolation view synthesis. Although these methods leverage the 3D data priors, they are limited by the scarcity and diversity of 3D data. Consequently, these methods struggle to acquire high-quality renderings in unseen areas, especially when OOD data is used as input.

### v2

@@ Abstract
Advancements in 3D scene reconstruction have transformed 2D images from the real world into 3D models, producing realistic 3D results from hundreds of input photos. Despite great success in dense-view reconstruction scenarios, rendering a detailed scene from sparse views is still an ill-posed optimization problem, often resulting in artifacts and distortions in unseen areas. In this paper, we propose ReconX, a novel 3D scene reconstruction paradigm that reframes the ambiguous reconstruction problem as a temporal generation task. The key insight is to unleash the strong generative prior of large pre-trained video diffusion models for sparse-view reconstruction. Nevertheless, it is challenging to preserve 3D view consistency when directly generating video frames from pre-trained models. To address this issue, given limited input views, the proposed ReconX first constructs a global point cloud and encodes it into a contextual space as the 3D structure condition. Guided by the condition, the video diffusion model then synthesizes video frames that are detail-preserved and exhibit a high degree of 3D consistency, ensuring the coherence of the scene from various perspectives. Finally, we recover the 3D scene from the generated video through a confidence-aware 3D Gaussian Splatting optimization scheme. Extensive experiments on various real-world datasets show the superiority of ReconX over state-of-the-art methods in terms of quality and generalizability. Project Page: https://liuff19.github.io/ReconX.
@@ 1 Introduction
To address the limitations, we propose ReconX, a novel 3D scene reconstruction paradigm that reformulates the inherently ambiguous reconstruction problem as a generation problem. Our key insight is to unleash the strong generative prior of pre-trained large video diffusion models (Blattmann et al., 2023a; b; Xing et al., 2023) to create more observations for the downstream reconstruction task. Despite the capability to synthesize video clips featuring plausible 3D structures (Gao et al., 2024), recovering a high-quality 3D scene from current video diffusion models is still challenging, due to the poor 3D view consistency across generated 2D frames. Grounded by theoretical analysis, we explore the potential of incorporating 3D structure condition into the video generative process, which bridges the gap between the under-determined 3D creation problem and the fully-observed 3D reconstruction setting. Specifically, given sparse images, we first build a global point cloud through a pose-free stereo reconstruction method. Then we encode it into a rich context representation space as the 3D condition in cross-attention layers, which guides the video diffusion model to synthesize detail-preserved frames with 3D consistent novel observations of the scene. Finally, we reconstruct the 3D scene from the generated video through Gaussian Splatting with a 3D confidence-aware and robust scene optimization scheme, which further deblurs the uncertainty in video frames effectively. Extensive experiments verify the efficacy of our framework and show that ReconX outperforms existing methods for high quality and generalizability, revealing the great potential to craft intricate 3D worlds from video diffusion models. The overview and examples of reconstructions are shown in Figure 1.
@@
In summary, our main contributions are as follows:
@@
  • We incorporate the 3D structure condition into the conditional space of the video diffusion model to generate 3D consistent frames and propose a 3D confidence-aware optimization scheme in 3DGS to reconstruct the scene given the generated video.
@@ 2 Related Work
Sparse-view reconstruction. NeRF and 3DGS typically demand hundreds of input images and rely on the multi-view stereo reconstruction (MVS) approach (e.g., COLMAP (Schönberger & Frahm, 2016)) to estimate the camera parameters. To address the issue of low-quality 3D reconstruction caused by sparse views, PixelNeRF (Yu et al., 2021) proposes using convolutional neural networks to extract features from the input context. Moreover, FreeNeRF (Yang et al., 2023) adopts the frequency and density regularized strategies to alleviate the artifacts caused by insufficient inputs without any additional cost. To mitigate the overfitting to input sparse views in 3DGS, FSGS (Zhu et al., 2023) and SparseGS (Xiong et al., 2023) employ a depth estimator to regularize the optimization process. However, these methods all require known camera intrinsics and extrinsics, which is not practical in real-world scenario. Benefiting from the existing powerful 3D reconstruction model (i.e., DUSt3R (Wang et al., 2024a)), InstantSplat (Fan et al., 2024) is able to acquire accurate camera parameters and initial 3D representations from unposed sparse-view inputs, leading to the efficient and high-quality 3D reconstruction.
@@
Regression model for generalizable view synthesis. While NeRF and 3DGS are optimized per-scene, a line of research aims to train feed-forward models that output a 3D representation directly from a few input images, bypassing the need for time-consuming optimization. Splatter image (Szymanowicz et al., 2024b) performs an efficient feed-forward manner for monocular 3D object reconstruction by predicting a 3D Gaussian for each image pixel. Meanwhile, pixelSplat (Charatan et al., 2024) proposes predicting the scene-level 3DGS from the image pairs, using the epipolar transformer to better extract scene features. Following that, MVsplat (Chen et al., 2024a) introduces the cost volume and depth refinements to produce a clean and high-quality 3D Gaussians in a faster way. LatentSplat (Wewer et al., 2024) encodes the variational 3D Gaussians and utilizes a discriminator to synthesize more realistic images. To reconstruct a complete scene from a single image, Flash3D (Szymanowicz et al., 2024a) adopts a hierarchical 3DGS learning policy and depth constraint to achieve high-quality interpolation and extrapolation view synthesis. Although these methods leverage the 3D data priors, they are limited by the scarcity and diversity of 3D data. Consequently, these methods struggle to achieve high-quality renderings in unseen areas, especially when out-of-distribution (OOD) data is used as input.